40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3847     Entry time: Tue Nov 2 16:24:07 2010
Author: Koji 
Type: Update 
Category: Auxiliary locking 
Subject: Alignment for the green in the X trans table  

[Kiwamu Koji]

Today we found the green beam from the end was totally missing at the vertex.

- What we found was very weak green beam at the end. Unhappy.

- We removed the PBS. We should obtain the beam for the fiber from the rejection of the (sort of) dichroic separator although the given space is not large.

- The temperature controller was off. We turned it on again.

- We found everything was still misaligned. Aligned the crystal, aligned the Faraday for the green.

- Aligned the last two steering mirrors such that we hit the approximate center of the ETMX and the center of the ITMX.

- Made the fine alignment to have the green beam at the PSL table.

The green beam emerged from the chamber looks not so round as there is a clipping at an in-vac steering.
We will make the thorough realignment before closing the tank.

ELOG V3.1.3-